If 118 people attend a concert and tickets for adults cost $3.75 while tickets for children cost $2.50 and total receipts for the concert was $366.25, how many of each went to the concert?
There were adults and children that attended the concert.

Answer by ikleyn(52873)

.
If 118 people attend a concert and tickets for adults cost $3.75
while tickets for children cost $2.50 and total receipts for the concert was $366.25,
how many of each went to the concert?
There were adults and children that attended the concert.
~~~~~~~~~~~~~~~~~~

If the number of adults is "a", then the number of children is (118-n).

With it, write the total money equation in the form

    3.75a + 2.50*(118-a) = 366.25  dollars.


Simplify and find "a"

    3.75a + 2.50*118 - 2.50a = 366.25

    3.75a - 2.50a = 366.25 - 2.50*118

        1.25a     =      71.25

            a     =      71/25/1.25 = 57.


ANSWER.  57 adults and 118-57 = 61 children.


CHECK.  57*3.75 + 61*2.50 = 366.25 dollars, total money.  ! correct !

Solved.

-----------------

To see many other similar problems solved,  look into the lesson
    - Typical problems on buying and selling items
in this site.